« Project Status: 7 Feb 2007 | Main | Status: 9FEB07 (Friday - sunrise) »

Project Status: 8 Feb 2007

Weather Status: IOP 4 came to an end at 9:20GMT on 8 Feb. Rainfall amounts almost made an inch (0.99 at Blue Canyon), but for the most part amounts over the ARB were less than that. Still, it was a good opportunity to get the systems up and running again. SMART-R ran flawlessly through the event. NSSL put together a nice radar loop documenting the event. IOP 5 start looks good with the sounding releases at Slough House and Reno beginning at 20 GMT/8 Feb. Precip should start sometime between 8/21GMT and 9/00GMT. The approaching trough does have a IPW river in association but amounts are less than expected from yesterday's models. Still the event will produce 3-5 inches of rain with the maximum rate near 9/18GMT. SAC has issued a winter storm warning for the mountains above 6000ft. There is a short break indicated in some forecast models between 10/03 and 10/18GMT as the first minor wave moves east and the onset of the major trough. Cold frontal passage with the major trough looks to be sometime near 11 Feb/06GMT but timing is not exact. Plan is to run sondes one cycle past the cold frontal passage. This will be updated at tomorrow's telecon. This system should continue to Sunday afternoon (11 Feb/21GMT). Next event isn't until 15 Feb.

System Status: SMART-R running well; First sondes have been released for IOP 5. Rain sampling is underway at Cazadero and Bodega Bay. Power outage at Big Bend has started and should be completed by late afternoon. Problems were encountered with ensemble products getting to the ALPS workstation at MRY; there appears to be a problem getting the high resolution model products using MS Explorer version 7.0. Earlier versions (6.x) appear to work. Tim Schneider will host the media for site visits and HMT interviews.

Operational Decision Summary: Radar for IOP 5 begins at operator discretion; Sondes are underway at RNO and Slough House. They will continue until at least Saturday night/Sunday morning (11 Feb/10GMT). Stay tuned for an update at tomorrows HMT telecon.

John McGinley ESRL/GSD

TrackBack

TrackBack URL for this entry:
/cgi-bin/mt/mt-tb.cgi/229